OpenStudy (anonymous):

8-x3 for x=1 -9 729 9 11 and the same thing goes its not x3 the 3 is small

OpenStudy (earthcitizen):

\[8-x^3\]

OpenStudy (earthcitizen):

nt quite clear

OpenStudy (anonymous):

neither am i

OpenStudy (earthcitizen):

can you rest8 the function ?

OpenStudy (anonymous):

i dont know sorry i dont get the question

OpenStudy (anonymous):

when you say that the 3 is small do you mean \[x^3\]

OpenStudy (anonymous):

oh yeah sorry

OpenStudy (anonymous):

do you know how to calculate \[(-9)^3\] ?

OpenStudy (anonymous):

kinda it depends haha sorry

OpenStudy (anonymous):

-9 x -9 = 81 81 * -9 = -729 8-(-729) = 8+729 = 737

OpenStudy (anonymous):

that doesnt make sence to me. sorry if i am making you mad or something

OpenStudy (anonymous):

how do you do the 3 in the small form?

OpenStudy (anonymous):

not at all a problem. when face with formula like this. consider the order of operations, the x3 happens first so for x= -9 x3 = -9 x -9 x -9 = 729

OpenStudy (anonymous):

oh i get it thanks(: